OpenStudy (anonymous):

solve by substitution or elimination: x^2+y^2=25 and x^2-y^2=7

hero (hero):

add them together

OpenStudy (anonymous):

\[x ^{2}+y ^{2}=25\] \[x ^{2}-y ^{2}=7\] Add both equation we get \[2x ^{2}=32\] implies \[x ^{2}=16\] x=+4 or -4

OpenStudy (anonymous):

they want ordered pairs as solutions.

OpenStudy (anonymous):

they want you to plug positive 4 back into one of the original equations and solve. same with the -4.

OpenStudy (anonymous):

what would that come out to be because im not getting the correct numbers?

OpenStudy (anonymous):

oh sorry i didn't completed the work

OpenStudy (anonymous):

Solution (x,y)=(4,3),(-4,3),(4,-3)

OpenStudy (whpalmer4):

Using the second equation: \[4^2-y^2=7\]\[16-7=y^2\]\[y=\pm\sqrt{9} = \pm3\] \[(-4^2-y^2=7)\]\[16-7=y^2\]\[y=\pm3\] Solutions are (4,3),(-4,3),(4,-3),(-4,-3)

OpenStudy (whpalmer4):

Sorry, I didn't get the parens in the right place, that should be \[(-4)^2-y^2=7\]but the outcome is the same.

OpenStudy (anonymous):

mr.whpalmer4 i forget to write (-4,-3)

OpenStudy (anonymous):

I'm confused about how it is 4 solutions, because plugging it in you don't get those. I'm not getting them.

OpenStudy (anonymous):

if \[x ^{2}=a\] then the solutions are +a or -a

OpenStudy (anonymous):

isn't the answer just + or - 4, and + or - 3?

OpenStudy (anonymous):

without writing the ordered pairs? isn't it the same thing?

OpenStudy (anonymous):

ordered pair used is to ensure the order

OpenStudy (anonymous):

That makes more sense. thank you so much Joseph, whpalmer, and Hero :)